# Personalized Item Embeddings in Federated Multimodal Recommendation

### v1

@@ Abstract.
Federated recommendation systems play a crucial role in protecting user privacy. However, existing methods primarily rely on ID-based item embeddings, overlooking the rich multimodal information of items. To address this limitation, we propose a novel Federated Multimodal Recommendation System called FedMR. FedMR leverages a foundation model on the server side to encode multimodal data, such as images and text, associated with items. To tackle the challenge of data heterogeneity caused by varying user preferences, FedMR introduces a Mixing Feature Fusion Module on the client. This module dynamically adjusts the weights of different fusion strategies based on user interaction history, generating personalized item embeddings that capture fine-grained user preferences. FedMR is compatible with existing ID-based federated recommendation systems, improving their performances without modifying the original framework. Our experiments on four real-world multimodal recommendation datasets demonstrate the effectiveness of FedMR. Our code is available at https://anonymous.4open.science/r/FedMR.
@@ 1. Introduction
Under privacy protection regulations, such as GDPR (Voigt and Von dem Bussche, 2017), that emphasize user data privacy, Federated Recommendation Systems (FedRec) have become an essential technology in modern life (Sun et al., 2024). However, in these systems, since each client typically only has access to a single user’s interaction history, data sparsity is a common issue that degrades recommendation performance (Li and Long, 2024). Modeling item features as shared information across clients can effectively mitigate the impact of data sparsity. Many studies (Zhang et al., 2023c; Li et al., 2023a, 2024b) have shown that personalized item feature representations can significantly improve recommendation results. Despite items often containing rich multimodal information (Zhou et al., 2023b), such as such as text and images, most existing FedRec methods (Zhang et al., 2023d, 2024d, 2024f) still rely mainly on ID embeddings, underutilizing the potential of multimodal data. Thus, generating personalized item embeddings using multimodal data while preserving user privacy remains a significant challenge.
@@
To address these challenges, Federated Multimodal Recommendation Systems (FedMMRec) have emerged, which integrates MMRec and introduces the capability of processing multimodal information into FedRec. By utilizing multimodal data and user-specific information, such as profiles and interaction history, locally on each client to generate personalized recommendations and aggregating model parameters at the server, FedMMRec enhances recommendation performance while preserving user privacy (Feng et al., 2024). FedMMRec is still in the early stages of research, and related studies are limited. FedMMR (Wu et al., 2024) adopts a model-splitting strategy where the computationally intensive components are handled by the server, while clients perform lightweight collaborative filtering. AMMFRS (Feng et al., 2024) introduces an attention mechanism on the client side to perform weighted fusion of multimodal data, effectively distinguishing and utilizing the contributions of different modalities. These methods not only protect user privacy but also effectively integrate multimodal features to improve recommendation accuracy and personalization. However, they rely on fixed multimodal fusion strategies, neglecting the personalized representation of item features on clients, which limits their ability to adapt to diverse user needs. Although AMMFRS uses an attention mechanism to distinguish the importance of different modalities, it lacks adaptability to the diversity of user data, resulting in limited flexibility and accuracy of fusion strategies (Niu et al., 2021).
Main Contribution. To integrate multimodal item information into FedRec while ensuring user privacy, we propose a Federated Multimodal Recommendation system (FedMR) based on foundation models. FedMR uses a foundation model on the server side to encode multimodal item data (e.g., images, text) into feature embeddings, and constructs a global item ID embedding to capture structural features among items. On each client, FedMR introduces a Mixing Feature Fusion Module that dynamically adjusts fusion strategies based on user interaction history, combining multimodal and ID embeddings to generate personalized item embeddings. This approach provides more personalized recommendations. Thus, FedMR seamlessly integrates multimodal information with existing ID-based FedRec methods, improving the flexibility and accuracy of recommendations without altering the existing framework, and effectively handles data heterogeneity across clients.
@@
  • A Mixing Feature Fusion Module is designed to dynamically adjusts fusion strategies based on user interaction histories to generate personalized item embeddings, enhancing personalization and flexibility in recommendations;
@@ 2.1. Federated Recommendation Systems
Federated recommendation systems (FedRec) aim to provide personalized recommendations while ensuring user privacy and data security (Yang et al., 2020). In recent years, FedRec has made significant progress in areas such as rating prediction (Lin et al., 2020a, b; Liang et al., 2021; Lin et al., 2022; Zhang et al., 2023d; Li et al., 2024a), top-k recommendation (Ammad-Ud-Din et al., 2019; Perifanis and Efraimidis, 2022; Luo et al., 2022; Zhang et al., 2023b; Dong et al., 2022; Li et al., 2022a; Qu et al., 2023), and addressing the cold-start problem (Liu et al., 2021b; Meihan et al., 2022; Wahab et al., 2022; Zhang et al., 2024c, d). However, most existing FedRec methods (Huang et al., 2020; Flanagan et al., 2021; Wu et al., 2022; Zhu et al., 2022; Imran et al., 2023; Yuan et al., 2023a, 2024) rely heavily on ID embeddings to model items, overlooking the rich multimodal information (e.g., titles and covers) that items may have. AMMFRS (Feng et al., 2024) and FedMMR (Wu et al., 2024) apply multimodal learning (Baltrušaitis et al., 2018) to FedRec. AMMFRS uses an attention mechanism (Niu et al., 2021) to weight and fuse multimodal data, emphasizing the contribution of different features. In contrast, FedMMR assigns the computationally heavy multimodal learning tasks to the server and introduces a secure cross-client convolution strategy to capture high-order interaction information using secret sharing techniques while preserving privacy. Unlike methods that focus only on personalizing user information, PFedRec (Zhang et al., 2023c) proposes a dual personalization mechanism that optimizes personalization for both users and items, integrating it into existing FedRec frameworks. Building on this, FedRAP (Li et al., 2023a) introduces additive personalization, which refines the modeling of user and item features. FedDAE (Li et al., 2024b) employs a dual-encoder variational autoencoder to adaptively fuse shared and personalized item knowledge through a personalized gating network. While these methods achieve personalized item embeddings, they still mainly rely on ID embeddings, failing to fully utilize the rich multimodal information of items.
@@ 4.2. Mixing Feature Fusion Module
The router $\mathbf{r}_{u}$ dynamically generates a weight vector $\mathbf{w}_{u}\in\mathbb{R}^{|\mathcal{G}|}$ based on the feature embeddings of items the user has interacted with, obtained through different fusion strategies, and assigns weights to each fusion strategy, i.e., the weight of the strategy $g_{j}$ is $w_{u,j}$. Then, we have the final fused features $\bar{\mathbf{F}}$ as follows:
(6) $\bar{\mathbf{F}}=\sum_{j=1}^{|\mathcal{G}|}w_{u,j}\mathbf{F}_{j}.$
According to Eq. (6), FedMR dynamically integrates and weights each user’s $u$ multimodal information and ID features to generate the personalized item embeddings $\bar{\mathbf{F}}$, enhancing the system’s ability to capture complex user interaction patterns.
@@ 4.3. Objective of FedMR
(7) $\min_{\{\theta_{u}\},\{\varphi_{u}\},\{\gamma_{j}\},\mathbf{D}}\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(f(\bar{\mathbf{F}};\theta_{u}),r_{u,i}).$
@@ 4.4. Framework
Upon receiving the feature data $\mathbf{V}$, $\mathbf{C}$ and $\mathbf{D}$ from the server, the client filters the embeddings of items that the user has interacted with and performs fusion using the MFFM. Within this module, a learnable user-specific router $\phi_{u}$ dynamically assigns weights to the fused embeddings for further weighted fusion. The fusion strategies include a Sum strategy, a MLP (Taud and Mas, 2018) strategy, and a Gate strategy. The fused features $\bar{\mathbf{F}}$ are then fed into the prediction function $f$ to generate predicted scores $\hat{\mathbf{r}}_{u}$ for items, and the local model parameters are optimized by minimizing reconstruction loss.
@@ 6. Conclusion
This paper presents FedMR, a federated multimodal recommendation system that incorporates a mixing feature fusion module on each client. This module dynamically adjusts fusion weights based on user interaction history, generating personalized item embeddings that capture fine-grained user preferences, thereby enhancing personalized recommendations. FedMR allows flexible modification of fusion strategies and supports different modalities of item data as needed. Experimental results show that FedMR performs well on four multimodal recommendation datasets and is compatible with existing ID-based methods without requiring changes to the underlying framework. This study offers new directions for the development of federated recommendation systems, particularly in the application of multimodal data and foundation models.

### v2

# Personalized Item Representations in Federated Multimodal Recommendation

@@ Abstract.
Federated recommendation systems are essential for providing personalized recommendations while protecting user privacy. However, current methods mainly rely on ID-based item embeddings, neglecting the rich multimodal information of items. To address this, we propose a Federated Multimodal Recommendation System, called FedMR. FedMR uses a foundation model on the server to encode multimodal item data, such as images and text. To handle data heterogeneity caused by user preference differences, FedMR introduces a Mixing Feature Fusion Module on each client, which adjusts fusion strategy weights based on user interaction history to generate personalized item representations that capture users’ fine-grained preferences. FedMR is compatible with existing ID-based federated recommendation systems, improving performance without modifying the original framework. Experiments on four real-world multimodal datasets demonstrate FedMR’s effectiveness. The code is available at https://anonymous.4open.science/r/FedMR.
@@ 1. Introduction
Under privacy protection regulations, such as GDPR (Voigt and Von dem Bussche, 2017), that emphasize user data privacy, Federated Recommendation Systems (FedRec) have become an essential technology in modern life (Sun et al., 2024). However, in these systems, since each client typically only has access to a single user’s interaction history, data sparsity is a common issue that degrades recommendation performance (Li and Long, 2024). Modeling item features as shared information across clients can effectively mitigate the impact of data sparsity. Many studies (Zhang et al., 2023c; Li et al., 2023a, 2024b) have shown that personalized item representations can significantly improve recommendation results. Despite items often containing rich multimodal information (Zhou et al., 2023b), such as such as text and images, most existing FedRec methods (Zhang et al., 2023d, 2024d, 2024f) still rely mainly on ID embeddings, underutilizing the potential of multimodal data. Thus, generating personalized item representations from multimodal data while ensuring user privacy remains a significant challenge.
@@
To address these challenges, Federated Multimodal Recommendation Systems (FedMMRec) have emerged, which integrates MMRec and introduces the capability of processing multimodal information into FedRec. By utilizing multimodal data and user-specific information, such as profiles and interaction history, locally on each client to generate personalized recommendations and aggregating model parameters at the server, FedMMRec enhances recommendation performance while preserving user privacy (Feng et al., 2024). To the best of our knowledge, FedMMRec is still in the early stages of research, and related studies are limited. FedMMR (Wu et al., 2024) adopts a model-splitting strategy where the computationally intensive components are handled by the server, while clients perform lightweight collaborative filtering. AMMFRS (Feng et al., 2024) introduces an attention mechanism on the client side to perform weighted fusion of multimodal data, effectively distinguishing and utilizing the contributions of different modalities. These methods not only protect user privacy but also effectively integrate multimodal features to improve recommendation accuracy and personalization. However, they rely on fixed multimodal fusion strategies, neglecting the personalized representation of item features on clients, which limits their ability to adapt to diverse user needs. Although AMMFRS uses an attention mechanism to distinguish the importance of different modalities, it lacks adaptability to the diversity of user data, resulting in limited flexibility and accuracy of fusion strategies (Niu et al., 2021).
Main Contribution. To integrate multimodal item information into FedRec while ensuring user privacy, we propose a novel Federated Multimodal Recommendation system (FedMR). FedMR uses a foundation model on the server to encode multimodal item data (e.g., images, text) into feature embeddings, and constructs a global item ID embedding to capture structural features among items. On each client, FedMR introduces a Mixing Feature Fusion Module that dynamically adjusts the importances of fusion strategies based on user interaction history, combining multimodal and ID embeddings to model fine-grained personalized item representations for users. Thus, FedMR seamlessly integrates multimodal information with existing ID-based FedRec methods, improving the flexibility and accuracy of recommendations without altering the existing framework, and effectively handles data heterogeneity across clients, providing more personalized recommendations.
@@
  • A Mixing Feature Fusion Module is designed to dynamically adjusts fusion strategies based on user interaction histories to generate personalized item representations, enhancing personalization and flexibility in recommendations;
@@ 2.1. Federated Recommendation Systems
Federated recommendation systems (FedRec) aim to provide personalized recommendations while ensuring user privacy and data security (Yang et al., 2020). In recent years, FedRec has made significant progress in areas such as rating prediction (Lin et al., 2020a, b; Liang et al., 2021; Lin et al., 2022; Zhang et al., 2023d; Li et al., 2024a), top-k recommendation (Ammad-Ud-Din et al., 2019; Perifanis and Efraimidis, 2022; Luo et al., 2022; Zhang et al., 2023b; Dong et al., 2022; Li et al., 2022a; Qu et al., 2023), and addressing the cold-start problem (Liu et al., 2021b; Meihan et al., 2022; Wahab et al., 2022; Zhang et al., 2024c, d). However, most existing FedRec methods (Huang et al., 2020; Flanagan et al., 2021; Wu et al., 2022; Zhu et al., 2022; Imran et al., 2023; Yuan et al., 2023a, 2024) rely solely on ID embeddings to model item representations, overlooking the rich multimodal information (e.g., titles and covers) that items may have. AMMFRS (Feng et al., 2024) and FedMMR (Wu et al., 2024) apply multimodal learning (Baltrušaitis et al., 2018) to FedRec. AMMFRS uses an attention mechanism (Niu et al., 2021) to weight and fuse multimodal data, emphasizing the contribution of different features. In contrast, FedMMR assigns the computationally heavy multimodal learning tasks to the server and introduces a secure cross-client convolution strategy to capture high-order interaction information using secret sharing techniques while preserving privacy. Unlike methods that focus only on personalizing user information, PFedRec (Zhang et al., 2023c) proposes a dual personalization mechanism that optimizes personalization for both users and items, integrating it into existing FedRec frameworks. Building on this, FedRAP (Li et al., 2023a) introduces additive personalization, which refines the modeling of user and item features. FedDAE (Li et al., 2024b) employs a dual-encoder variational autoencoder to adaptively fuse shared and personalized item knowledge through a personalized gating network. While these methods achieve personalized item representations, they still mainly rely on ID embeddings, failing to fully utilize the rich multimodal information of items.
@@ 4.2. Mixing Feature Fusion Module
The router $\mathbf{r}_{u}$ dynamically generates a weight vector $\mathbf{w}_{u}\in\mathbb{R}^{|\mathcal{G}|}$ based on the feature embeddings of items the user has interacted with, obtained through different fusion strategies, and assigns weights to each fusion strategy, i.e., the weight of the strategy $g_{j}$ is $w_{u,j}$. Then, we have the final fused features $\bar{\mathbf{F}}^{(u)}$ as follows:
(6) $\bar{\mathbf{F}}^{(u)}=\sum_{j=1}^{|\mathcal{G}|}w_{u,j}\mathbf{F}_{j}.$
According to Eq. (6), FedMR dynamically integrates and weights each user’s $u$ multimodal information and ID features to generate the personalized item representations $\bar{\mathbf{F}}^{(u)}$, thereby enhancing the system’s ability to capture fine-grained user preferences.
@@ 4.3. Objective of FedMR
(7) $\min_{\{\theta_{u}\},\{\varphi_{u}\},\{\gamma_{j}\},\mathbf{D}}\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(f(\bar{\mathbf{F}}^{(u)};\theta_{u}),r_{u,i}).$
@@ 4.4. Framework
Upon receiving the features $\mathbf{V}$, $\mathbf{C}$ and $\mathbf{D}$ from server, the client filters the embeddings of items that the user has interacted with and performs fusion using the MFFM. Within this module, a learnable user-specific router $\phi_{u}$ dynamically assigns weights to the fused embeddings for further weighted fusion. The fusion strategies include a Sum strategy, a MLP (Taud and Mas, 2018) strategy, and a Gate strategy. The personalized item representations $\bar{\mathbf{F}}^{(u)}$ are then fed into a prediction function $f$ to generate predicted scores $\hat{\mathbf{r}}_{u}$, and optimize the model parameters by minimizing reconstruction loss in Eq. (7).
@@ 6. Conclusion
This paper presents FedMR, a federated multimodal recommendation system that incorporates a mixing feature fusion module on each client. This module dynamically adjusts fusion weights based on user interaction history, generating personalized item representations that capture fine-grained user preferences, thereby enhancing personalized recommendations. It allows flexible modification of fusion strategies and supports different modalities of item data as needed. Experimental results on four multimodal recommendation datasets show that FedMR performs well and is compatible with existing ID-based methods without requiring changes to the underlying framework. This study offers new directions for the development of federated recommendation systems, particularly in the application of multimodal data and foundation models.